3)For the following reaction, Kc = 255 at 1000 K.
CO (g) + Cl2 (g) ⇌ COCl2 (g)

Part A
If a reaction mixture initially contains a CO concentration of 0.1530 and a Cl2 concentration of 0.174 at 1000K. What is the equilibrium concentration of CO at 1000 K?
Part B
What is the equilibrium concentration of Cl2 at 1000 K?
Part C
What is the equilibrium concentration of COCl2 at 1000 K?

4)Consider the following reaction:
HC2H3O2(aq)+H2O(l)⇌H3O+(aq)+C2H3O−2(aq)
Kc=1.8×10−5 at 25∘C

If a solution initially contains 0.180 M HC2H3O2, what is the equilibrium concentration of H3O+ at 25∘C?

5)Consider the following reaction:
Br2(g)+Cl2(g)⇌2BrCl(g)
Kp=1.11×10−4 at 150 K.
A reaction mixture initially contains a Br2 partial pressure of 755 torr and a Cl2 partial pressure of 740 torr at 150 K.

Calculate the equilibrium partial pressure of BrCl.
1
Expert's answer
2020-03-30T10:53:11-0400

Phosgene dissociates almost completely at this temperature, so the applied equilibrium constant for the reverse reaction

COCl2=CO+Cl2

[CO][Cl 2 ] / [COCl 2 ]=255 mole/ l

[COCl2]=X moles

[CO]=0,153-x=0.1527958

[Cl 2 ]=0,174-x=0.173895

=>(0.153-x) (0.174-x)/x=255

0.026622-0.327 x+x2=255x

x=0.0001042


[H3O]+[C2HO2]-/[(HCOOH)2]

Let x = change in [HC2H3O2]

At equilibrium: [HC2H3O2] = (0.250 – x)

At equilibrium = [H3O+] = x and [C2H3O2-] = x

1.8 x 10^-5 = (x) (x) / (0.250 – x) ; I will neglect x in the (0.180 –x) since x is much smaller than 0.250

x^2 = 1.8 x 10^-5*(0.180)

x^2 = 1.8*1.8*10-6

x = [H3O+] = 1.8*10-3